OpenStudy (anonymous):

How much energy is contained in 4.2 x 106 photons with a wavelength of 92 μm?

OpenStudy (anonymous):

First, use the formula, E=hc/lambda, where E is the energy for 1photon, h is Planck's constant or 6.626*10^-34 Joules and c is the speed of light or 3*10^8 m/s. Make lambda subject of the formula. Can you do this and tell me what you get?

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

Yes but I don't don't think I'm setting this up right, I am stumped

OpenStudy (anonymous):

Could you tell me how to send an attachment?

OpenStudy (anonymous):

I usually just take a picture with my phone and put it onto my computer and send it as an attachment....but if you have like paint or something you could do it in there, save it and attach that

OpenStudy (anonymous):

Hmm, I cant... Anyway, to simplify , you can first multiply h snd c, because they are constants. You will get 1.99*10^-25. E is the energy for 1 photon, then 4.2*10^6 have 4.2*10^6E. So your new formula is E=1.99*10^-25/92*10^-6. You simplify it to E=2.16*10^-21. This is the energy for a photon, but since you have 4.2*10^6, you multiplu the E with this number, getting 9.072*10^-15J. Got it?

OpenStudy (anonymous):

Yes I do, thank you!

OpenStudy (anonymous):

You are welcome.